OpenStudy (anonymous):

What is the equivalent resistance as an algebraic expression (in terms of R) of network A as viewed from its port? incorrect

OpenStudy (anonymous):

3R or (R+R+R)

OpenStudy (nurali):

@manimuthu Given that each individual resistor can dissipate up to 1 watt of power before burning up, how much total power in watts ( W ) can the smallest-valued composite resistor dissipate before burning up? P=V^2/Rmin_______(1) P=1 Watt Rmin=R1=R2=4 Ohms, and R3= 6 Ohms Find V In equation 1 1=V^2/4 V=2 Put V in equation 1 P=V^2/R3 P=(2)^2/6 P=2.667 W

OpenStudy (nurali):

if correct please click best response.